# The Price of Competitive Information Disclosure

### v1

@@ 2.2 Individual Utilities and Social Welfare
[…]
@@
Denote the agents selected without tying with others by $\mathcal{W}(\bm{\sigma})$, and agents tying with others by $\mathcal{T}(\bm{\sigma})$. We have
@@ 3 Lower Bounds on the Price of Anarchy
In this section, we establish lower bounds on the price of anarchy (PoA) for our problem. We do so by explicitly evaluating $\mathrm{PoA}(\mathcal{I})$ under a simple setting, where the principalselects only one agent (i.e., $k=1$), and with $N$ symmetric agents with Bernoulli values—where for every $i\in[N]$, $\mathcal{F}_{i}$ is an i.i.d. $\mathtt{Bernoulli}(\zeta)$ distribution—and constant utility of selection—where all agents get utility $u_{i}(0)=u_{i}(1)=1$ for $v\in\{0,1\}$ when selected. This setting is considered by Au and Kawai (2020), who show that in this case, a symmetric Nash equilibrium exists, and is unique.
@@
[…]
@@
Let $\hat{p}=\zeta\cdot N\leq 1$ by assumption. We first note that the first-best social welfare for this setting is the probability at least one agent has value $1$, i.e.,
@@
Turning to signaling, observe that if two signals $\sigma_{i,1}$ and $\sigma_{i,2}$ share the same posterior mean, combining these into a single signal affects neither agent’s utilities nor welfare. Therefore, without loss of generality, we can assume that there is at most one signal on each posterior mean value. For convenience, we use a signal’s posterior mean to represent the signal, thus $\Sigma_{i}\subseteq[0,1]$ for all agents $i$, and for any $\sigma_{i}\in\Sigma_{i}$, it holds that $e_{i}(\sigma_{i})=\sigma_{i}$. The signaling scheme $\mathcal{Z}_{i}$ can therefore be represented by a posterior mean distribution (or equivalently, induced signal distribution) $\Pi_{i}\in\Delta([0,1])$.
For agent $i$, denote the cumulative distribution function (cdf) of strategy $\Pi_{i}$ as $G_{i}(v)=\Pr_{x\sim\Pi_{i}}[x\leq v]$ for $v\in[0,1]$. By the work of Au and Kawai (2020, Section 3.1), since $\hat{p}=\zeta\cdot N<1$, the symmetric Nash equilibrium (that is, all $N$ agents play the same strategy) exists, and in this equilibrium, each agent’s strategy $\Pi_{i}=\Pi$ has the same cumulative distribution $G_{i}=G$ as follows:
@@
Let $\rho_{i}(v;G)$ denote the probability that each agent $i$ is selected conditioned on sending a signal (with mean) $v$ and choosing the strategies of all other agents according to $G$. Then, it follows that
$\rho_{i}(v;G)=\left(G(v)\right)^{N-1}=\begin{cases}\frac{v}{\hat{p}}&\quad\text{when }p\in[0,\hat{p}],\\ 1&\quad\text{when }p\in(\hat{p},1].\end{cases}$
@@
Since the equilibrium is symmetric, the overall social welfare is $\mathrm{SW}(\bm{\mathcal{Z}})=N\cdot\mathrm{SW}_{i}(\bm{\mathcal{Z}})=\frac{N\hat{p}}{2N-1}$. Therefore, combining with Eq. 1, we get
$\mathrm{PoA}(\mathcal{I})=\frac{\mathrm{SW}^{\max}(\mathcal{I})}{\mathrm{SW}(\bm{\mathcal{Z}})}\geq\frac{(2N-1)}{N\hat{p}}\cdot\left(1-e^{-\hat{p}}\right)=\left(2-\frac{1}{N}\right)\cdot\left(\frac{1-e^{-\hat{p}}}{\hat{p}}\right).$
This completes the proof of Theorem 3.1. ∎
By Theorem 3.1, when $\zeta$ is close to 00 (so that $\hat{p}=N\zeta$ is small), the lower bound on the price of anarchy approaches $2$.
@@ 4 Warm-up: Identical Priors and Constant Utility of Selection
In this section, as a warm-up, we present a constant upper bound on PoA when all agents have identical priors, each agent gets constant utility from being selected, and the number of agents selected is $k=1$. The ideas developed here form the basis for the more elaborate analysis on the case of general priors and utility functions in the next section. We show the following theorem.
@@
$E_{(q)}=\frac{\int_{1-q}^{1}v(x)\mathop{\mathrm{d}\!}x}{q}=\frac{1}{q}\cdot\int_{1-q}^{1}v(x)\mathop{\mathrm{d}\!}x.$
Note that by definition, $q\cdot E_{(q)}$ is non-decreasing in $q\in[0,1]$.
@@
Now consider any Nash Equilibrium $\bm{\mathcal{Z}}$, and let $M$ be the random variable representing the highest posterior mean among all agents. Since the principalselects the agent with the highest posterior mean, we have $\mathrm{SW}(\bm{\mathcal{Z}})=\mathbb{E}[M]$. The technical crux of our argument, formalized below in Lemma 4.2, lies in showing that $\Pr[M\geq E_{(2/N)}]\geq 1/2$. Combining these, we get
@@
This completes the proof of Theorem 4.1.
@@
Note that by construction of $\pi_{i^{*}}^{\prime}$, the posterior mean of agent $i^{*}$’s value under signal $s_{1}$ is $E_{(2/N)}$. Thus, whenever $M_{-i^{*}}<E_{(2/N)}$ and agent $i^{*}$ sends $s_{1}$, she is selected by the principal. By Eq. 2, conditioned on sending signal $s_{1}$, agent $i^{*}$ is selected with probability strictly greater than $1/2$. Moreover, since $s_{1}$ is sent with probability $2/N$, agent $i^{*}$’s expected utility by deviating to $\mathcal{Z}_{i^{*}}^{\prime}$ is at least
Pr[s1 is $\Pr[\text{$s_{1}$ is sent}]\cdot\Pr[\text{$i^{*}$ is selected}\mid\text{$s_{1}$ is sent}]>\frac{1}{2}\cdot\frac{2}{N}=\frac{1}{N}.$
@@ 4.2 Discussion
Key ideas.
A central idea in our analysis is the use of quantile cuts to capture the upper tail of each agent’s value. In the warm-up case, since all agents share the same prior and $k=1$, we apply a uniform $\frac{1}{N}$ quantile cut to each agent’s distribution. This directly yields an upper bound on the maximum social welfare as the conditional mean of the values in the top $\frac{1}{N}$ quantile. Even when agents have non-identical priors, we can define a personalized quantile for each agent. Though the cut may vary across the agents, the core idea remains the same: We focus on the upper tail to obtain an upper bound of the optimal social welfare.
@@
Challenges with generalization.
In the warm-up case, all agents share the same prior $\mathcal{F}$. This symmetry allows us to define a single quantile function $v(q)$ and we can easily upper bound the social welfare by using a uniform $\frac{1}{N}$ quantile cut on each agent’s prior. However, when agents’ priors are non-i.i.d., the optimal selection may involve different thresholds for different agents, and we need to analyze their contributions to the social welfare using heterogeneous quantile cuts.
When the principalselects more than one agent (i.e., $k>1$), the analysis becomes more challenging, since different agents should be treated differently based on their distributions, and the corresponding quantile cuts, which could be different even for comparable distributions.
@@ 5 Constant Factor Upper Bound on PoA
The PoA of the Bayesian persuasion game with $k$ selection (for any $k\geq 1$) is at most $11+5\sqrt{5}$ when the priors $\mathcal{F}_{i}$ are arbitrary independent distributions, and the functions $u_{i}(\cdot)$ are arbitrary monotone.
@@
For simplicity, we will assume that we are working on a given instance $\mathcal{I}$ and omit the dependence on $\mathcal{I}$ in our notations in this section. Our analysis will involve parameters $\alpha,\beta,\tau,\phi$ and $Q$ that we will optimize later. The parameters must satisfy the following constraints:
$\alpha,\beta,\tau,\phi,Q>0,\qquad\tau<1,\qquad Q\leq 1,\qquad\alpha\cdot Q>1,\qquad\beta<\tau,\qquad\phi\cdot\tau\geq\alpha.$
@@ 5.1 Top Quantile Cuts
For any $q\in[0,1]$, let $v_{i}(q)$ denote the $q$-th quantile of $\mathcal{F}_{i}$. For simplicity, we assume that $\mathcal{F}_{i}$ contains no point masses, and we will address in footnotes when the argument is different when point masses occur. Our first step is to perform a line sweep of the distributions $\mathcal{F}_{i}$ so that the total quantiles in the sweep sum up to $k\cdot Q$. To do so, we first choose a value $E_{\mathrm{cut}}$, and define agent $i$’s quantile cut
@@
In other words, $q_{i}$ is the largest fraction $x$ such that the conditional mean of the top $x$ quantile is at least $E_{\text{cut}}$. The threshold $E_{\text{cut}}$ is chosen to satisfy the total quantile cut constraint: [^8] Note that if $E_{\mathrm{cut}}$ equals the highest value of some $\mathcal{F}_{i}$ and $\mathcal{F}_{i}$ has a point mass at that value, satisfying Eq. 3 may become impossible. A slight increase in $E_{\mathrm{cut}}$ can cause the corresponding agents’ $q_{i}$ to drop abruptly to 00. In such cases, we define $E_{\mathrm{cut}}$ as $\sup\{x\in[0,1]:\sum_{i\in[N]}q_{i}\geq k\cdot Q\}$ and then reduce those $q_{i}$ values on the largest value mass until $\sum_{i\in[N]}q_{i}=k\cdot Q$.
$\sum_{i\in[N]}q_{i}=k\cdot Q.$ (3)
For agent $i$, denote by $E_{i}=\mathbb{E}[v_{i}(q)\mid q>1-q_{i}]=\frac{\int_{1-q_{i}}^{1}v_{i}(q)\mathop{\mathrm{d}\!}q}{q_{i}}$ her conditioned value mean above the quantile cut $q_{i}$. [^9] If $q_{i}=0$, we simply set $E_{i}=E_{\mathrm{cut}}$. By definition, we have $E_{i}\geq E_{\mathrm{cut}}$ for all $i\in[N]$. We must note that there may be agents for whom $q_{i}=1$. For such agents, their conditional mean above $q_{i}$ (i.e., $E_{i}$) can differ from the common threshold $E_{\mathrm{cut}}$, because the line stops moving before the sweep process ends. See Fig. 1 for an illustration.
Figure 1: An illustration of the final state of the line sweep for an example with $5$ agents. As shown in the figure, agent $2$ and agent $5$ have a mean above the sweep line larger than $E_{\mathrm{cut}}$, as their quantile cuts satisfy $q_{i}=1$.
@@
Since in this scenario the principalalways selects a (weak) superset of the agents selected under full-revelation of values (where the principalknows the revealed value of all agents and selects the top $k$ posterior mean agents), the social welfare achieved under full revelation of values $\mathrm{SW}^{\max}$ is at most $\mathrm{SW}^{\prime}$.
@@
Fix a constant $\alpha>1/Q$. Since $Q\leq 1$, we have $\alpha>1$. We group the agents into two sets $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ based on their $r_{i}(\bm{\mathcal{Z}})$ as:
@@
Since $\sum_{i\in[N]}r_{i}(\bm{\mathcal{Z}})\leq k$, we have $\sum_{i\in\mathcal{N}_{1}}q_{i}\leq k/\alpha$. Therefore, $\sum_{i\in\mathcal{N}_{2}}q_{i}\geq(Q-1/\alpha)\cdot k$. Since $Q>1/\alpha$, $\mathcal{N}_{2}$ cannot be an empty set. We have the following simple lemma.
@@
In the rest of the analysis, we will fix a constant $\beta>0$. We split the analysis based on the value generated by agents in $\mathcal{N}_{2}$ when they are selected.
5.2 Case 1: All Agents in $\mathcal{N}_{2}$ are Large Contributors
For each agent $i$, define the value contribution of $i$ as

### v2

@@ 2.2 Individual Utilities and Social Welfare
If agent $i\in[N]$ is selected with value $v$, she receives utility $u_{i}(v)$. We assume that the function $u_{i}\colon[0,1]\to\mathbb{R}_{\geq 0}$ is continuous and monotonically increasing, with the property that $u_{i}(v)>0$ for any v>0v>0.999Otherwise, if $u_{i}(v)\equiv 0$ for all agents, the PoA can become trivially unbounded, since any strategy profile will be a Nash equilibrium.
@@
Denote the agents selected without tying with others by $\mathcal{W}(\bm{\sigma})$, and denote agent(s) with posterior mean equal to $e_{(k)}(\bm{\sigma})$ (i.e., those who possibly form a tie and get selected) by $\mathcal{T}(\bm{\sigma})$. We have
@@ 3 Lower Bounds on the Price of Anarchy
In this section, we establish lower bounds on the price of anarchy (PoA) for our problem. We do so by explicitly evaluating $\mathrm{PoA}(\mathcal{I})$ under a simple setting, where the principal selects only one agent (i.e., $k=1$), and with $N$ symmetric agents with Bernoulli values—where for every $\in[N], i\mathcal{F}_{i}$ is an i.i.d. $\mathtt{Bernoulli}(\zeta)$ distribution—and constant utility of selection—where all agents get utility $u_{i}(0)=u_{i}(1)=1$ for $v\in\{0,1\}$ when selected. This setting is considered by Au and Kawai (2020), who show that in this case, a symmetric Nash equilibrium exists, and is unique.
@@
$\mathrm{PoA}(\mathcal{I})\geq\left(2-\frac{1}{N}\right)\left(\frac{1-e^{-N\zeta}}{N\zeta}\right)\in\left[1.264,\ 2-O\left(\frac{1}{N}\right)\right].$
@@
Let $\hat{p}=\zeta\cdot N\leq 1$ by assumption. We first note that the first-best social welfare for this setting is the probability that at least one agent has value $1$, i.e.,
@@
Turning to signaling, observe that if two signals $\sigma_{i,1}$ and $\sigma_{i,2}$ share the same posterior mean, combining these into a single signal affects neither agent $i$’s utility nor welfare. Therefore, without loss of generality, we can assume that there is at most one signal for each posterior mean value. For convenience, we use a signal’s posterior mean to represent the signal, thus $\Sigma_{i}\subseteq[0,1]$ for all agents $i$, and for any $\sigma_{i}\in\Sigma_{i}$, it holds that $e_{i}(\sigma_{i})=\sigma_{i}$. The signaling scheme $\mathcal{Z}_{i}$ can therefore be represented by a posterior mean distribution (or equivalently, induced signal distribution) $\Pi_{i}\in\Delta([0,1])$.
For agent $i$, denote the cumulative distribution function (cdf) of strategy $\Pi_{i}$ by $G_{i}(v)=\Pr_{x\sim\Pi_{i}}[x\leq v]$ for $v\in[0,1]$. By the work of Au and Kawai (2020, Section 3.1), since $\hat{p}=\zeta\cdot N<1$, the symmetric Nash equilibrium (that is, all $N$ agents play the same strategy) exists, and in this equilibrium, each agent’s strategy $\Pi_{i}=\Pi$ has the same cumulative distribution $G_{i}=G$ as follows:
@@
Let $\rho_{i}(v;G)$ denote the probability that agent $i$ is selected conditioned on sending a signal (with mean) $v$ and choosing the strategies of all other agents according to $G$. Then, it follows that
$\rho_{i}(v;G)=\left(G(v)\right)^{N-1}=\begin{cases}\frac{v}{\hat{p}}&\quad\text{when }v\in[0,\hat{p}],\\ 1&\quad\text{when }v\in(\hat{p},1].\end{cases}$
@@
Since the equilibrium is symmetric, the overall social welfare is $\mathrm{SW}(\bm{\mathcal{Z}})=N\cdot\mathrm{SW}_{i}(\bm{\mathcal{Z}})=\frac{N\hat{p}}{2N-1}$. Therefore, combining with Eq.˜1, we get
[…]
This completes the proof of Theorem˜3.1. ∎
By Theorem˜3.1, when $\zeta$ is close to 0 (so that $\hat{p}=N\zeta$ is small), the lower bound on the price of anarchy approaches $2$.
Unbounded PoA for negative-valued agents.
We can construct an example where the PoA is unbounded when agents’ values can be negative from the lower bound example above. Suppose we shift every agent’s value distribution to the left by $\mathrm{SW}_{i}(\mathcal{Z})-\varepsilon$. Every agent using the same signaling scheme remains in equilibrium since their utilities only depend on whether they are selected. Therefore, the expected welfare in the equilibrium is $\varepsilon$. The $\mathrm{SW}_{\max}$ is also decreased by the same gap amount, but will be much larger than $\varepsilon$ (since originally there is a constant gap between these two welfare values). This leads to an unbounded PoA when the values can be negative.
@@ 4 Warm-up: Identical Priors and Constant Utility of Selection
In this section, as a warm-up, we present a constant upper bound on PoA when all agents have identical priors, each agent gets constant utility from being selected, and the number of agents selected is $k=1$. The ideas developed here form the basis for the more elaborate analysis of general priors and utility functions in the next section. We show the following theorem.
@@
$E_{(q)}=\frac{1}{q}\cdot\int_{1-q}^{1}v(x)\mathop{\mathrm{d}\!}x.$
Note that by definition, $q\cdot E_{(q)}$ is non-decreasing in $q\in(0,1]$.
@@
Now consider any Nash Equilibrium $\bm{\mathcal{Z}}$, and let $M$ be the random variable representing the highest posterior mean among all agents. Since the principal selects the agent with the highest posterior mean, we have $\mathrm{SW}(\bm{\mathcal{Z}})=\mathbb{E}[M]$. The technical crux of our argument, formalized below in Lemma˜4.2, lies in showing that $\Pr[M\geq E_{(2/N)}]\geq 1/2$. Combining these, we get
@@
This completes the proof of Theorem˜4.1.
@@
Note that by construction of $\pi_{i^{*}}^{\prime}$, the posterior mean of agent $i^{*}$’s value under signal $s_{1}$ is $E_{(2/N)}$. Thus, whenever $M_{-i^{*}}<E_{(2/N)}$ and agent $i^{*}$ sends $s_{1}$, she is selected by the principal. By Eq.˜2, conditioned on sending signal $s_{1}$, agent $i^{*}$ is selected with probability strictly greater than $1/2$. Moreover, since $s_{1}$ is sent with probability $2/N$, agent $i^{*}$’s expected utility by deviating to $\mathcal{Z}_{i^{*}}^{\prime}$ is at least
$\Pr[\text{$s_{1}$ is sent}]\cdot\Pr[\text{$i^{*}$ is selected}\mid\text{$s_{1}$ is sent}]>\frac{1}{2}\cdot\frac{2}{N}=\frac{1}{N}.$
@@ 4.2 Discussion
A central idea in our analysis is the use of quantile cuts to capture the upper tail of each agent’s value. In the warm-up case, since all agents share the same prior and $k=1$, we apply a uniform $\frac{1}{N}$ quantile cut to each agent’s distribution. This directly yields an upper bound on the maximum social welfare as the conditional mean of the values in the top $\frac{1}{N}$ quantile. When agents have non-identical priors, a natural idea is that we can define a personalized quantile for each agent. Though the cut may vary across the agents, the core idea remains the same: We focus on the upper tail to obtain an upper bound of the optimal social welfare.
However, using a similar approach requires more elaboration. Since now we must deal with heterogeneity between agents and $k$ can be larger than $1$, it may be impossible to construct an equal-mean cut among all agents like the warm-up case. We need to carefully handle the edge-case agents whose values are too large compared to others since the sweeping line may exhaust their entire distribution. Moreover, this method also makes the social welfare cap a combination of these heterogeneous tails, which adds more challenges to our analysis.
@@ 5 Constant Factor Upper Bound on PoA
The PoA of the Bayesian persuasion game with $k$ selection (for any $k\geq 1$) is at most $11+5\sqrt{5}$ when the priors $\mathcal{F}_{i}$ are arbitrary independent distributions, and the functions $u_{i}(\cdot)$ are positive and weakly monotone.
@@
For simplicity, we will assume that we are working on a given instance $\mathcal{I}$ and omit the dependence on $\mathcal{I}$ in our notations in this section. Our analysis will involve parameters $\alpha,\beta,\tau$ and $\phi$ that we will optimize later. The parameters must satisfy the following constraints:
$\alpha,\beta,\tau,\phi>0,\qquad\tau<1,\qquad\alpha>1,\qquad\beta<\tau,\qquad\phi\cdot\tau\geq\alpha.$
@@ 5.1 Top Quantile Cuts
For any $q\in[0,1]$, let $v_{i}(q)$ denote the $q$-th quantile of $\mathcal{F}_{i}$. For simplicity, we assume that $\mathcal{F}_{i}$ contains no point masses, and we will address in footnotes when the argument is different when point masses occur. Our first step is to perform a line sweep of the distributions $\mathcal{F}_{i}$ so that the total quantiles in the sweep sum up to $k$. To do so, we first choose a value $E_{\mathrm{cut}}$, and define agent $i$’s quantile cut
@@
In other words, $q_{i}$ is the largest fraction $x$ such that the conditional mean of the top $x$ quantile is at least $E_{\text{cut}}$. The threshold $E_{\text{cut}}$ is chosen to satisfy the total quantile cut constraint: [^12] Note that if $E_{\mathrm{cut}}$ equals the highest value of some $\mathcal{F}_{i}$ and $\mathcal{F}_{i}$ has a point mass at that value, satisfying Eq. 3 may become impossible. A slight increase in $E_{\mathrm{cut}}$ can cause the corresponding agent’s $q_{i}$ to drop abruptly to 0. In such cases, we define $E_{\mathrm{cut}}$ as $\sup\{x\in[0,1]:\sum_{i\in[N]}q_{i}\geq k\}$ and then reduce those $q_{i}$ values on the largest value mass until $\sum_{i\in[N]}q_{i}=k$.
$\sum_{i\in[N]}q_{i}=k.$ (3)
For agent $i$, denote by $E_{i}=\mathbb{E}[v_{i}(q)\mid q>1-q_{i}]=\frac{\int_{1-q_{i}}^{1}v_{i}(q)\mathop{\mathrm{d}\!}q}{q_{i}}$ her conditioned value mean above the quantile cut $q_{i}$. [^13] If $q_{i}=0$, we simply set $E_{i}=E_{\mathrm{cut}}$. By definition, we have $E_{i}\geq E_{\mathrm{cut}}$ for all $i\in[N]$. We must note that there may be agents for whom $q_{i}=1$. For such agents, their conditional mean above $q_{i}$ (i.e., $E_{i}$) can differ from the common threshold $E_{\mathrm{cut}}$, because the line stops moving before the sweep process ends. See Fig.˜1 for an illustration.
Figure 1: An illustration of the final state of the line sweep for an example with 55 agents. As shown in the figure, agent $2$ and agent 55 have a mean above the sweep line larger than $E_{\mathrm{cut}}$, as their quantile cuts satisfy $q_{i}=1$.
@@
Since in this scenario the principal always selects a (weak) superset of the agents selected under full-revelation of values (where the principal knows the revealed value of all agents and selects the $k$ agents with top values), the social welfare achieved under full revelation of values $\mathrm{SW}^{\max}$ is at most $\mathrm{SW}^{\prime}$.
@@
Fix a constant $\alpha>1$. We group the agents into two sets $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ based on their $r_{i}(\bm{\mathcal{Z}})$ as:
@@
Since $\sum_{i\in[N]}r_{i}(\bm{\mathcal{Z}})\leq k$, we have $\sum_{i\in\mathcal{N}_{1}}q_{i}\leq k/\alpha$. Therefore, $\sum_{i\in\mathcal{N}_{2}}q_{i}\geq(1-1/\alpha)\cdot k$. Since $1/\alpha<1$, $\mathcal{N}_{2}$ cannot be an empty set. We have the following simple lemma.
@@
In the rest of the analysis, we will fix a constant $\beta>0$. We split the analysis based on the value generated by agents in $\mathcal{N}_{2}$ when they are selected. For each agent $i$, define the value contribution of $i$ as